OpenStudy (anonymous):

Find x, if the point (x, −12) is on the terminal side of θ and sin θ = − 4/5.

OpenStudy (dumbcow):

 using SOH CAH TOA \[\sin \theta = \frac{-12}{r} = \frac{-4}{5}\] solve for "r" then use pythagorean thm to get "x"

OpenStudy (anonymous):

Wow, thanks man! ^^ I appreciate the guidance.